A photographer wishes to mix 5 liters of 4% acetic acid solution with a 8% solution to get a 7% solution. How many liters of 8% solution must be added?
Answer by stanbon(75887)

A photographer wishes to mix 5 liters of 4% acetic acid solution with a 8% solution to get a 7% solution. How many liters of 8% solution must be added?
---
Equation:
acid + acid = acid
0.04*5 + 0.08x = 0.07(5+x)
-----
Multiply thru by 100 to get:
4*5 + 8x = 7*5 + 7*x
x = 3*5
x = 15 liters (amt. of 8% solution needed)
==============================================
Cheers,
Stan H.
===========